Let f be a linear function. If f(−3) = 7and f(2) = −13,find f(x).
Answer by scott8148(6628)

linear means ___ f(x) = mx + b

7 = m(-3) + b ___ 7 = -3m + b

-13 = m(2) + b ___ -13 = 2m + b

subtracting equations (eliminating b) ___ 20 = -5m ___ -4 = m

substituting ___ 7 = -3(-4) + b ___ -5 = b

f(x) = -4x - 5